The length of a rectangle is 1cm greater than twice its width. The area of the rectangle is 120cm^2.
Find the length and width.

Answer by Alan3354(69443)

The length of a rectangle is 1cm greater than twice its width. The area of the rectangle is 120cm^2.
Find the length and width.
---------------
Area = L*W = 120
L = 2W + 1
W*(2W + 1) = 120
2W^2 + W -120 = 0
(2W - 15)*(W + 8) = 0
W = 7.5
L = 16